How would the equation for the following problem be written?
The length of a rectangle is 8 cm more than 3 times its width. The perimeter of the rectangle is 64 cm.

Answer by algebrahouse.com(1659)

"How would the equation for the following problem be written?
The length of a rectangle is 8 cm more than 3 times its width. The perimeter of the rectangle is 64 cm."

x = width
3x + 8 = length {length is 8 more than 3 times width}

Perimeter of a rectangle = (2 x width) + (2 x length)

2x + 2(3x + 8) = 64 {perimeter = 2(width) + 2(length)}
2x + 6x + 16 = 64 {used distributive property}
8x + 16 = 64 {combined like terms}
8x = 48 {subtracted 16 from both sides}
x = 6 {divided both sides by 8}
3x + 8 = 26 {substituted 6, in for x, into 3x + 8}

width = 6 and length = 26
www.algebrahouse.com